a rope 36 feet long is cut into three pieces, the second piece is four feet longer tha the first, the las t piece is three times as long as the second. if x represents the length of the first piece, then which equation determines the lenght of the first piece?
36=5x+8
36=x+(x+4)+(3x)
36=3x+12
36=x+(x+4)+3(x+4)
36=3x+16

Answer by oberobic(2304)

A rope is cut into three pieces: x, y, and z.
y = x+4
z = 3(x+4
.
substitute
.
x + (x+4) + 3(x+4) = 36
.
x + x + 4 + 3x + 12 = 36
.
5x + 16 = 36
.
5x = 20
.
x = 4
.
y = 4+4 = 8
.
z = 3(8) = 24
.
4 + 8 + 24 = 36
.
Correct.
.
Done.